# FreqDGT: Frequency-Adaptive Dynamic Graph Networks with Transformer for Cross-subject EEG Emotion Recognition

### v2

@@ 3.2 Adaptive Dynamic Graph Learning
where $\mathbf{D}_{dyn}[i,i]=\sum_{j}\tilde{\mathbf{A}}_{dyn}[i,j]$ represents the degree matrix.
@@ 4.1 Datasets and Experimental Protocol
FAP Components w/o Cross-band Attention 79.3% 60.1% 69.2%
@@
ADGL Components w/o Fixed Adjacency 78.5% 59.4% 68.3%
@@
MTDN Components w/o Multi-scale Transformer 79.1% 60.2% 69.4%
@@ Acknowledgments
This work was supported by The Hong Kong Polytechnic University Start-up Fund (Project ID: P0053210), the Hong Kong Polytechnic University Faculty Reserve Fund (Project ID: P0053738), an internal grant from The Hong Kong Polytechnic University (Project ID: P0048377), The Hong Kong Polytechnic University Departmental Collaborative Research Fund (Project ID: P0056428) and The Hong Kong Polytechnic University Collaborative Research with World-leading Research Groups Fund (Project ID: P0058097).

### v3

@@ 3.2 Adaptive Dynamic Graph Learning
where $\mathbf{D}_{dyn}[i,i]=\sum_{j}\tilde{\mathbf{A}}_{dyn}[i,j]$ represents the degree matrix. Multi-level graph convolution with $K$-th order Chebyshev approximation propagates features:
$\mathbf{H}^{(l+1)}=\sum_{k=0}^{K-1}\theta_{k}^{(l)}T_{k}(\tilde{\mathbf{L}})\mathbf{H}^{(l)}$ (11)
[…]
@@ 4.1 Datasets and Experimental Protocol
FAP
Components
w/o Cross-band Attention 79.3% 60.1% 69.2%
@@
ADGL
Components
w/o Fixed Adjacency 78.5% 59.4% 68.3%
@@
MTDN
Components
w/o Multi-scale Transformer 79.1% 60.2% 69.4%
@@ Acknowledgments
This work was supported by The Hong Kong Polytechnic University Start-up Fund (Project ID: P0053210), the Hong Kong Polytechnic University Faculty Reserve Fund (Project ID: P0053738), an internal grant from The Hong Kong Polytechnic University (Project ID: P0048377), The Hong Kong Polytechnic University Departmental Collaborative Research Fund (Project ID: P0056428), The Hong Kong Polytechnic University Collaborative Research with World-leading Research Groups Fund (Project ID: P0058097) and The Research Grants Council Collaborative Research Fund (Project ID: P0049774).